# Differentially Private Next-Token Prediction of Large Language Models

### v1

@@ 9 Ethical Considerations
Additionally, we plan to publicly release our code under the Apache License, Version 2.0. We believe that allowing open access to these experiments will help spark academic research of this work, as well as protect the privacy of user data in commercial deployment of large language models.
@@ Appendix A Additional Experimental Details
Appendix C Properties of RDP
@@
$\epsilon^{{}^{\prime}}_{q}(\alpha)=\frac{1}{\alpha-1}\log\left((1-q)^{\alpha-1}(1+(\alpha-1)q)\\ +\sum_{k=2}^{\alpha}\binom{\alpha}{k}(1-q)^{\alpha-k}q^{k}e^{(k-1)\epsilon(k)}\right).$ (12)
@@ Appendix D Supplementary Experimental Figures
Figure 5 shows the utility trade-off of $\alpha$. Hence, it is crucial to the performance of PMixED that $\alpha$ is small.

### v3

@@ 9 Ethical Considerations
Additionally, we publicly released our code under the Apache License, Version 2.0. We believe that allowing open access to these experiments will help spark academic research of this work, as well as protect the privacy of user data in commercial deployment of large language models.
Acknowledgements
The authors would like to thank Tony Ginart for his tremendous advice and guidance early on in the project. This work is supported by Defense Advanced Research Projects Agency (DARPA) under Contract Nos. HR001120C0088, NSF award number 2224319, DGE-1842487, REAL@USC-Meta center, and gifts from Google and VMware. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.
@@ Appendix A Additional Experimental Details
Appendix C Properties of Renyi Divergence and RDP
@@
{dmath}
ϵ^’_q(α) = 1α-1log((1-q)^α-1 (1 + (α- 1)q)
+ ∑_k=2^α (αk) (1-q)^α- k q^k e^(k-1) ϵ(k) ).
@@
Theorem C.7 (Convexity in Second Argument Van Erven and Harremos (2014)).
For any order $\alpha\in[0,\infty]$ Renyi divergence is convex in its second argument. That is, for any probability distributions $P,Q,Q^{\prime}$ and $s\in[0,1]$
$D_{\alpha}(P||(1-s)Q^{\prime}+sQ^{\prime})\leq(1-s)D_{\alpha}(P||Q^{\prime})+D_{\alpha}(P||Q).$
@@ Appendix D Supplementary Experimental Figures
Figure 5 shows the utility trade-off of $\alpha$. Hence, it is crucial to the performance of PMixED that $\alpha$ is small due to the monotonicity property of Renyi Divergence.
Appendix E Extended Related Works
Another promising privacy-related notion, machine unlearning, has emerged to reduce memorization by verifiably removing learned information from a data sample without retraining a model from scratch Guo et al. (2019); Bourtoule et al. (2021). Generally speaking, there are two machine unlearning notions: (1) exact unlearning, where the resulting model has completely unlearned a data sample Bourtoule et al. (2021), and (2) approximate unlearning, where a data point has been unlearned to some degree with high probability. It is well-known that differential privacy implies approximate machine unlearning. One recent, orthogonal work explored the use of task vectors to perform memorization unlearning of the private dataset Gao et al. (2024).